# Accurate Estimation of Mutual Information in High Dimensional Data

### v1

@@ 4.2 Estimator Performance with Finite Datasets
In the infinite-data regime, neural MI estimators receive a fresh data batch at every training step, which mitigates overfitting. Yet, in practice, datasets are always finite, which may change the behavior of estimators. The impact of a finite sample size on the estimators remains poorly understood [31]; here we address this gap.
The Max-Test Heuristic.
Figure 3 demonstrates overfitting for finite data in the high-dimensional teacher model of Fig. 2 (ground-truth MI of 4 bits). We track $I_{\rm InfoNCE}$ and $I_{\rm SMILE}$ on training and held-out sets for $2562^{8}=$ 256 (undersampled) and $3842^{14}=16,$ 384 (better sampled) pairs of $K=500$-dimensional $X$ and $Y$ as training progresses. Because 64 epochs in the undersampled setting expose the network to the same number of examples as a single epoch in the better-sampled case (but with repetition), direct epoch counts are not comparable. In both regimes, the training curves start below the true MI and then rise (InfoNCE until its usual saturation) as the networks fit finer-scale structure of the critics. Because the training curves show no clear inflection when they surpass the ground-truth, common heuristics (fixed epoch counts or loss plateau) give no reliable stopping signal. In contrast, the test curves grow initially but soon collapse, revealing overfitting.
To devise a better stopping rule, note that a neural critic gradually refines a smoothed approximation to the DV-optimal critic, and training time effectively sets the smallest resolved scale. Early stopping oversmooths and underestimates MI, whereas late stopping undersmooths, placing a spike at each training point, thus pushing the training MI high while the test MI falls. This mirrors kernel density estimation (KDE) of MI [52]. The optimal resolution must depend on data complexity and sample size, so no fixed epoch rule can work universally.
@@
Figure 3: The max-test heuristic. We evaluate neural MI estimators in a finite-data setting using the teacher model from Fig. 2, where 10 latent variables carrying 4 bits of MI are embedded in 500-dimensional $X$ and $Y$. We compare two sampling regimes for InfoNCE (left) and SMILE (right): 256 samples (under-sampled) and a larger dataset of $3842^{14}=16{,}$ 384 samples (better-sampled). In all cases, the test-set MI initially rises before declining due to overfitting (we do not show the negative values). The max-test heuristic selects the epoch with the peak test MI and uses the corresponding training MI as the reported estimate. Here the batch size is 128, so that InfoNCE is not in danger of saturating.
@@
In KDE-based MI estimation, one chooses the bandwidth that maximizes held-out likelihood [16]. Since the joint-density term dominates test error, this is the same as maximizing test MI. Analogously, we introduce the max-test heuristic: track the MI estimate on a test batch each epoch, pick the epoch where this test estimate peaks, and report the corresponding training MI as the final value (see Appx. A.3 for implementation details). To our knowledge this stopping strategy, which we use in all subsequent experiments, has not been formalized for neural MI estimation.
Although we report the training MI at the max-test epoch, this is not ad-hoc. This explicitly selects the critic with the best generalization. However, for MI estimation, generalization, as estimated via the test loss, is not the goal per se. Instead, one aims to approximate a functional of the joint distribution. The critic learns this functional on the training data, and this estimate—rather than the held-out test result—should be reported (see Appx. A.3). With finite sample size $N$, the max-test critic cannot capture arbitrarily fine structure of the data distribution. Hence, we view the max-test MI as an unbiased estimate of information at a certain resolution, set by $N$, and it may still underestimate the true MI. The degree of underestimation depends on the data, critic architecture, and the optimizer.
@@
Figure 4 compares InfoNCE, SMILE, and their probabilistic variants for varying sample sizes $N$ and for different ground truth MI for the $K=500$, $K_{Z}=10$ frozen random teacher networks model. For small MI (4 bits), all estimators converge to truth when $K_{Z}\ll N\allowbreak\sim O(10^{2})<K$. Thus MI estimation requires good sampling of the latent space, but not the full data space. As before, InfoNCE exhibits lower variance than SMILE.
[…]
Low-dimensional Latent Structure Is Needed for Reliable Estimation.
[…]
@@
Figure 5: Effect of latent and critic dimensionality on InfoNCE. Curves show mean $\pm$ s.d. over ten runs. Panel represents $K=500$-dimensional data generated by teacher networks with latent dimensionality $K_{Z}=10,100,$ and $500$ (left to right). The true MI is 4 bits throughout. A sufficiently expressive critic ($k_{Z}\geq K_{Z}$) is required to recover all the information, yet the estimate approaches 4 bits only in the low-dimensional latent case ($K_{Z}\ll K$) when sample size satisfies $N\gg K_{Z}$. For larger latent spaces, the estimate remains far below the target even with large $N$. Vertical lines mark $N$ needed for detection of nonzero MI using Gaussian random matrix models in the latent space ($N^{*}_{Z}$) and in the full space ($N^{*}$), cf. Appx. A.5 ($N^{*}_{Z}\approx 1$ not shown in the left panel; and $N^{*}_{Z}=N^{*}$ in the right panel). Since a nonzero estimate emerges at $N>N^{*}_{Z}$, but $N<N^{*}$ if $K_{Z}\ll K$, sampling of the latent space (not the full data space) governs the estimation even in the non-Gaussian setting.
@@
Figure 5 shows results for InfoNCE; the other estimators behave similarly (not shown). As expected, when the latent space is small ($K_{Z}=10$, left), a critic with $k_{Z}\geq K_{Z}$ captures all dependencies once $N\gg K_{Z}$, and the MI estimate approaches the truth. With a moderate latent dimension ($K_{Z}=100$, center), setting $k_{Z}<K_{Z}$ still fails outright, while $k_{Z}\geq K_{Z}$ retrieves only part of the information: even $N\sim 10^{5}$ samples leave the estimate well below four bits. In the extreme case where the latent and observed dimensions coincide ($K_{Z}=K=500$, right), neither an over-sized critic nor $N\sim 10^{5}$ samples is sufficient to recover meaningful information. The rising latent dimensionality inflates the sample requirement in two ways: it increases the effective dimension, for which generic methods demand more data [25, 22, 31], and it spreads the total information over many weaker dependencies, each harder to detect without still larger datasets. Surprisingly, the number of samples needed to estimate nonzero MI matches signal detection limits in spiked covariance models from random matrix theory [53, 54] in the latent and not the data space, cf. Appx. A.5.
@@
Thus, accurate MI estimation requires that: (i) the dependencies can be represented in a low-dimensional latent space, (ii) the critic is expressive enough to model that space, and (iii) the dataset is sufficient to detect signals in that space, with the size rising quadratically with latent dimension (cf. Appx. A.5). Therefore, in high dimensions, compressive embeddings used by the estimators are not just advantageous but essential.
@@ 4.3 Practical Guide to MI Estimation
Because no estimator can be uniformly unbiased for all distributions [39], sample-size dependent bias is the dominant practical concern. The accepted remedy is empirical: test how the estimate changes when the sample size is varied near the actual dataset size [6, 55, 3]. The best practice protocol randomly partitions the data into $\gamma=1,2,3,\dots$ equal subsets and computes MI for each subset, $I_{\mu}(\gamma)$, $\mu=1,\dots,\gamma$. One then checks whether $I_{\mu}(\gamma)$ is statistically stable as $\gamma$ varies [40], or whether it extrapolates linearly to the hypothetical $\gamma=0$ limit (effectively, infinite data) [6]. Pronounced curvature in $I_{\mu}(\gamma)$ vs $\gamma$ signals that the estimator is not in its large $N$ asymptotic regime and is therefore unreliable. In addition, the scatter of $I_{\mu}(\gamma)$ at fixed $\gamma$ provides an empirical variance that can be projected to $\gamma=1$ (the full dataset) or to $\gamma=0$ (infinite data) [40].
We combine these best practices with our analysis above to suggest the following algorithm for neural MI estimation (see Appx. A.4 for details):
@@
Figure 6 illustrates the workflow on data from a random teacher model with $K=500$, $K_{Z}=10$ and true MI of 4 bits. The left and the right panels show the undersampled ($N=256$) and the well-sampled ($N=2^{14}$) cases, respectively. On the left, using $k^{*}_{Z}=32$, we can reliably extrapolate $I_{\text{InfoNCE}}$ to $N\to\infty$ limit ($\gamma=0$) via subsampling. On the right, the estimator is already stable at this $N$; its slight downward bias matches Fig. 5 (see Discussion). We stress how striking this result is: in a 500$\times$500-dimensional, highly nonlinear setting (but with a 10-d latent structure), we obtain a near-perfect MI estimate, complete with accurate error bars, from just 256 samples!
A Real-World Example: Noisy MNIST.
@@
We next run our pipeline on a realistic noisy MNIST dataset [56, 57, 58, 19]. Each sample consists of two $28\times 28=784$ dimensional views: $X$ is a randomly rotated and scaled image of a digit, and $Y$ is another random digit with the same label overlaid with Perlin noise (see Appx. Fig. 11). The only shared information is the digit label (10 classes), giving a ground truth MI of $\log_{2}10\approx 3.3$ bits. The observed dimension is high ($K=784$), while the latent dimension should be low (but not precisely known) because only ten classes matter. Random matrix analysis, Appx. A.5, suggests that the nonzero MI detection thresholds are $N^{*}_{Z}\approx 17$ (for $K_{Z}=10)$ or $N^{*}\approx 2700$. Numerical experiments show that $N\sim 512$ is sufficient for detection, again suggesting that good sampling in the latent space is more important than in the data space. Here, however, we are interested in accurate estimation of MI, rather than just significant difference from zero.
@@ 5 Discussion
Accurate MI estimation in high-dimensional data remains difficult because no estimator is uniformly unbiased. Neural estimators offer flexibility, but most studies test them in well-sampled, linear settings that hide their limitations. We show that neural estimators become reliable when three requirements hold: the statistical dependence lies in a low-dimensional latent space; the critic is expressive enough to capture that space; and the dataset is large enough to resolve it. Under these conditions, three simple additions make neural MI estimation practical. First, a “max-test” early-stopping rule prevents the runaway growth common in DV objectives. Second, using the MINE family critics within a probabilistic VSIB wrapper regularizes them at high information values. Third, a subsampling-and-extrapolation workflow detects bias, chooses the right critic dimension, and estimates information with error bars. Although we illustrated this with InfoNCE and specific network architectures, our analysis is agnostic to both, and any DV-based estimator can be treated similarly.
Using these ingredients we obtained near-exact MI estimates, with error bars, in regimes that defeat classical methods (cf. SI Fig. 10), e.g., 784-dimensional image pairs with only $\sim 10^{4}$ samples. Across all experiments, our procedure never significantly overshot the ground truth, a crucial property because overestimation produces false positives in typical scientific applications [6, 16, 3], whereas modest underestimation vanishes as $N$ grows and finer structure is learned. This is also why the small downward bias sometimes visible at large $N$ (cf. Fig. 6 (right)), which is an artifact of a fixed batch size unable to explore fine-scale data features, is not a major concern.

### v2

@@ 4.2 Estimator Performance with Finite Datasets
For infinite-data, neural estimators receive a fresh data batch at every training step. Yet, in practice, datasets are always finite. The impact of a finite sample size on the estimators remains poorly understood (though see Czyz et al. (2023)); here we address this gap.
@@
Figure 3: The stopping heuristic. We evaluate neural MI estimators for finite-data using the teacher model from Fig. 2, where 10 latent variables carrying 4 bits of MI are embedded in 500-dimensional $X$ and $Y$. We compare two sampling regimes for InfoNCE (left) and SMILE (right): 256 samples (under-sampled) and a larger dataset of $2^{14}=16{,}384$ samples (better-sampled). In all cases, the test-set MI initially rises before declining due to overfitting (we do not show the negative values). The stopping heuristic selects the epoch with the peak test MI but reports the corresponding training MI. Here the batch size is 128, so that InfoNCE does not saturate.
The Stopping Heuristic.
Figure 3 demonstrates overfitting for finite data in the high-dimensional teacher model of Fig. 2 (ground-truth MI of 4 bits). We track $I_{\rm InfoNCE}$ and $I_{\rm SMILE}$ on training and held‑out sets for $2^{8}=256$ (undersampled) and $2^{14}=16,384$ (better sampled) pairs of $K=500$-dimensional $X$ and $Y$ as training progresses. Because 64 epochs in the undersampled setting expose the network to the same number of examples as a single epoch in the better-sampled case (but with repetition), direct epoch counts are not comparable. All training curves start below the true MI and then rise (InfoNCE until the saturation) as the networks fit finer‑scale structure of the critics. Because the training curves show no clear inflection when they surpass the ground‑truth, common heuristics (fixed epoch counts or loss plateau) give no reliable stopping signal. In contrast, the test curves grow initially but soon collapse, revealing overfitting.
Figure 4: MI vs. sample size for low and high information. We compare InfoNCE, SMILE, and VSIB versions with the max-test stopping for different sample sizes. Data from the frozen teacher model ($10$ latent, $500$ data dimensions). All estimators use separable critics, $k_{z}=32$. Means $\pm$ s.d. over 10 trials shown. Left: For small MI (4 bits), all estimators recover the ground truth for $10^{2}\lesssim N<K=500$. Right: For high MI (8 bits), contrastive estimators (InfoNCE and $\text{InfoNCE}-\text{VSIB}$) saturate near $\log(\text{batch size})=7$ bits. SMILE overestimates dramatically as NN grows. $\text{SMILE}-\text{VSIB}$ tracks the ground truth accurately for all $N\gtrsim 10^{2}$.
To devise a better stopping rule, note that the training time effectively sets the smallest scale resolved in the neural approximation of the DV-optimal critic. Early stopping oversmooths and underestimates MI, whereas late stopping undersmooths, thus pushing the training MI high while the test MI falls. This mirrors kernel density estimation (KDE) of MI (Moon et al., 1995). The optimal resolution must depend on data complexity and sample size, so no fixed epoch rule would work.
@@
In KDE MI estimation, one chooses the bandwidth that maximizes held-out likelihood (Margolin et al., 2006). Since the joint-density term dominates test error, this maximizes test MI. Analogously, our stopping heuristic is: track MI on a test batch each epoch, pick the epoch where this test value peaks, and report the corresponding training MI as the estimate (see Appx. A.3 for justification). To our knowledge, such rule has not been formalized for neural MI estimation.
@@
Figure 4 compares InfoNCE, SMILE, and their probabilistic variants for varying sample sizes and different ground truth MI for the $K=500$, $K_{Z}=10$ random teacher networks model. For small MI (4 bits), all estimators converge to truth when $K_{Z}\ll N\allowbreak\sim O(10^{2})<K$. Thus MI estimation requires good sampling of the latent (but not the data) space. As before, InfoNCE has lower variance than SMILE.
For high true MI (8 bits), all variational bounds and corresponding estimators degrade (Poole et al., 2019). InfoNCE and $\text{InfoNCE}-\text{VSIB}$ both saturate near $\log(\text{batch size})=7$ bits, as expected (van den Oord et al., 2018). Consistent with Figs. 1, 2, SMILE substantially overestimates when large sample size allows overtraining. In contrast, $\text{SMILE}-\text{VSIB}$ remains accurate and stable, converging to the correct value at $K_{Z}\ll N\lesssim K_{X}+K_{Y}$. This highlights the utility of the new probabilistic critic family. By treating the critic output as a distribution (e.g., Gaussian) and regularizing via the encoding terms (Appx. A.2), these critics mitigate the pathological variance and overfitting in SMILE. Their benefit is most pronounced for high MI, where resolving fine-grained data structure is essential.
Low-dimensional Latent Structure Allows Reliable Estimation.
Our experiments have used data with statistical dependencies in a relatively low-dimensional latent space, $K_{Z}=10$. We now ask how the latent dimensionality limits neural MI estimation by simultaneously varying the true latent dimension, $K_{Z}$, and the critic’s embedding dimension, $k_{Z}$, while keeping the data dimension $K=500$ large and fixed. We consider a low-dimensional latent setting, $K_{Z}=1$, a moderate regime, $1\ll K_{Z}=100\ll K$, and a fully high-dimensional latent regime, $1\ll K_{Z}=K=500$. With the ground-truth MI at 4 bits (where all estimators can work, Fig. 4), each canonical pair contributes $0.4$, $0.04$, and $0.008$ bits (equivalent correlation $\rho\approx 0.65$, $0.23$, and $0.11$, respectively), before they are nonlinearly mixed by the teacher networks. We expect that, if $k_{Z}\geq K_{Z}$, the critics can recover dependencies in the low-dimensional case (similar to Abdelaleem et al. (2024)). However, as $K_{Z}$ increases, the critic must disentangle an ever-growing number of weak interactions, and its MI estimate will deteriorate at a fixed $N$.
@@
Figure 5: Effect of latent and critic dimensionality on InfoNCE. Curves show mean $\pm$ s.d. over ten runs. Panel represents $K=500$-dimensional data generated by teacher networks with latent dimensionality $K_{Z}=10,100,$ and $500$ (left to right). The true MI is 4 bits throughout. A sufficiently expressive critic ($k_{Z}\geq K_{Z}$) is required to recover all the information, yet the estimate approaches 4 bits only in the low-dimensional latent case ($K_{Z}\ll K$) when sample size satisfies $N\gg K_{Z}$. For larger latent spaces, the estimate remains far below the target even with large $N$. Vertical lines mark NN needed for detection of nonzero MI using Gaussian random matrix models in the latent space ($N^{*}_{Z}$) and in the full space ($N^{*}$), cf. Appx. A.5 ($N^{*}_{Z}\approx 1$ not shown in the left panel; and $N^{*}_{Z}=N^{*}$ in the right panel). Since a nonzero estimate emerges at $N>N^{*}_{Z}$, but $N<N^{*}$ if $K_{Z}\ll K$, sampling of the latent space (not the full data space) governs the estimation even in the non-Gaussian setting.
@@
Figure 5 shows results for InfoNCE; not shown estimators behave similarly. When the latent space is small ($K_{Z}=10$, left), a critic with $k_{Z}\geq K_{Z}$ captures all dependencies once $N\gg K_{Z}$, and the estimate approaches the truth. With a moderate latent dimension ($K_{Z}=100$, center), setting $k_{Z}<K_{Z}$ still fails outright, while $k_{Z}\geq K_{Z}$ retrieves only part of the information: even $N\sim 10^{5}$ samples result in <4<4 bits. If the latent and observed dimensions coincide ($K_{Z}=K=500$, right), neither an over-sized critic nor $N\sim 10^{5}$ samples is sufficient to recover information. The sample requirement in affected in two ways: for larger effective dimension, generic methods demand more data (Khan et al., 2007; Gao et al., 2015; Czyz et al., 2023), and contribution of each dimension to the total information is smaller and harder to detect. Surprisingly, the number of samples needed to estimate nonzero MI matches signal detection limits in spiked covariance models from random matrix theory (Baik et al., 2005; Potters & Bouchaud, 2020) in the latent and not the data space (Appx. A.5).
@@
Thus, accurate MI estimation requires that: (i) the dependencies can be represented in a low-dimensional latent space, (ii) the critic is expressive enough to model that space, and (iii) the dataset is sufficient to sample that space, with the size rising quadratically with latent dimension (Appx. A.5). Therefore, in high dimensions, compressive embeddings are not just advantageous but essential. Using methods such as in Kandasamy et al. (2015), it should be possible to formalize these results and show that, as long as sufficiently smooth embedding can be constructed, reliable estimation with fast convergence with the data set size should be possible directly in the embedding space.
@@ 4.3 Practical Guide to MI Estimation
Because no estimator is unbiased for all distributions (Paninski, 2003), sample-size dependent bias is the dominant practical problem. The remedy is empirical: test how the estimate changes when the sample size varies (Strong et al., 1998; Nemenman et al., 2004; Tang et al., 2014). The best practice randomly partitions the data into $\gamma=1,2,3,\dots$ equal subsets and computes MI for each subset, $I_{\mu}(\gamma)$, $\mu=1,\dots,\gamma$. One checks if $I_{\mu}(\gamma)$ is statistically stable as $\gamma$ varies (Holmes & Nemenman, 2019), or if it extrapolates linearly to the hypothetical $\gamma=0$ limit (effectively, infinite data) (Strong et al., 1998). Large curvature in $I_{\mu}(\gamma)$ vs $\gamma$ signals that the estimator is not in its asymptotic regime and is unreliable. In addition, the scatter of $I_{\mu}(\gamma)$ at fixed $\gamma$ provides an empirical variance that can be extrapolated to $\gamma=1$ (all data) or $\gamma=0$ (infinite data) (Holmes & Nemenman, 2019).
We combine these best practices with our analysis into the following guidelines for neural MI estimation (see Appx. A.4 for details):
@@
Figure 6 illustrates the workflow on data from a random teacher model with $K=500$, $K_{Z}=10$ and true MI of 4 bits. The left and the right panels show the undersampled ($N=256$) and the well-sampled ($N=2^{14}$) cases, respectively. On the left, using $k^{*}_{Z}=32$, we can reliably extrapolate $I_{\text{InfoNCE}}$ to $N\to\infty$ limit ($\gamma=0$) via subsampling. On the right, the estimator is already stable at this $N$; its slight downward bias matches Fig. 5 (see Discussion). We stress how striking this result is: in a $\times500$-dimensional, highly nonlinear setting (but with a 10-d latent structure), we obtain a near-perfect MI estimate, complete with accurate error bars, from just 256 samples!
Table 1: Comparison to selected benchmarks of Czyz et al. (2023) (results in nats to match the reference).
Task True MI simple InfoNCE Ours
Asinh @ Student-t 3 × 3 0.29 0.20 0.26 ± 0.03
Multinormal 25 × 25 1.29 1.20 1.26 ± 0.03
Multinormal 50 × 50 1.62 1.40 1.60 ± 0.03
Normal CDF @ Multinormal 25 × 25 1.02 0.80 0.94 ± 0.12
Spiral @ Multinormal 25 × 25 1.02 0.70 0.97 ± 0.24
Student-t 3 × 3 0.29 0.20 0.26 ± 0.04
Student-t 3 × 3 0.18 0.10 0.18 ± 0.03
Evaluations I: Benchmarking.
We illustrate advantages of our pipeline in a subset of standard benchmarks (Czyz et al., 2023) in Tbl. 1. Our approach consistently estimates MI better than other methods, while also providing error bars. Additional comprehensive benchmarks in Appx. A.7.3.
Evaluations II: A Real-World Example, Noisy MNIST.
We apply our pipeline to a realistic noisy MNIST dataset (LeCun et al., 1998; Wang et al., 2015; 2016; Abdelaleem et al., 2025). Each sample consists of two $28\times 28=784$ dimensional views: $X$ is a randomly rotated and scaled image of a digit, and $Y$ is another random digit with the same label overlaid with Perlin noise (Appx. Fig. 11). The only shared information is the digit label (10 classes), giving the ground truth MI of $\log_{2}10\approx 3.3$ bits. The data dimension is $K=784$, while the latent dimension is low (but unknown) because only ten classes matter. Random matrix analysis, Appx. A.5, suggests that the nonzero MI detection thresholds are $N^{*}_{Z}\approx 17$ (for $K_{Z}=10)$ or $N^{*}\approx 2700$. Numerical experiments show that $N\sim 512$ is sufficient for detection, so that, again, good sampling in the latent (not data) space is sufficient. Here, however, we need accurate estimation of MI, rather than just significant difference from zero.
@@ 5 Discussion
Accurate MI estimation for high-dimensional data remains difficult because no estimator is uniformly unbiased. Neural estimators offer flexibility, but most studies test them in well-sampled, linear settings that hide their limitations. We show that neural estimators become reliable when: the statistical dependence lies in a low-dimensional latent space; the critic is expressive enough to capture that space; and the dataset is large enough to resolve it. Under these conditions, three improvements make neural MI estimation practical. First, a “max-test” early-stopping rule prevents the runaway growth common in DV objectives. Second, using the MINE family critics within a probabilistic VSIB wrapper regularizes them at high information values. Third, a subsampling-and-extrapolation workflow detects bias, chooses the right critic dimension, and estimates error bars. Although we illustrated this with InfoNCE and specific network architectures, our analysis can incorporate any DV-based estimator.
Using these ingredients we obtained near-exact MI estimates, with error bars, in regimes that defeat most methods (Appx. Fig. 9), e.g., 784-dimensional images with only $\sim 10^{4}$ samples. Across all experiments, our procedure never significantly overshot the ground truth, which is crucial to not produce false positives in typical scientific applications (Strong et al., 1998; Margolin et al., 2006; Tang et al., 2014), whereas modest underestimation vanishes as $N$ grows and finer structure is learned. This is also why the small downward bias sometimes visible at large $N$ (cf. Fig. 6 (right)), which is an artifact of a fixed batch size unable to explore fine-scale data features, is not a major concern.